Find the average value gave of the function g on the given interval. g(x) = 5x2 1 + x3, [0, 2]
Answer #1

The average value of over an interval [a, b Fav = te s fresda function ] is given flae) by - we have function 1 glas= 5x² [it€ Also from h we have - When x=0, t= It 2013 t=1 & When He te Itajo t = 18 t = 9 -6 3 olor alus Using ②, ③, ④, ③ in 1 we getgav = ( 27-1] gav = 5 x 26 lav = 8 van dan Answel. .
Answer #2

SOLUTION :


Average value of the given function in interval 0 to 2 :


   2

= ∫  f(x) dx  / (2 - 0)

   0


            2

=   1/2 ∫  5x^2 sqrt(1 + x^3) dx

           0


Let 1 + x^3 = t

=> 3x^2 dx = dt

=> 5 x^2 dx = 5dt/3 


At x = 0, t = 1 and at x = 2, t = 9



So, average value :


          9                                                      9

= 1/2 ∫  5/3 sqrt(t) dt = 5/6 * 2/3 [ t^(3/2) ]                            

         1                                                       1


= 5/9 [9^(3/2) - 1^(3/2)]

= 5/9 * 26

= 130/9 = 14.44 (ANSWER)

answered by: Tulsiram Garg